OpenStudy (torii897):

Which of the following values in the set below will make the equation 2x + 1 = 3 true? {0, 1, 2, 3, 4}

OpenStudy (texaschic101):

just solve for x 2x + 1 = 3 --- subtract 1 from both sides 2x = 3 - 1 2x = 2 -- now divide both sides by 2 x = 1.....any other value for x will not be true

OpenStudy (torii897):

thank you wow ur good lol

OpenStudy (texaschic101):

you can even check it by subbing in 1 for x 2(1) + 1 = 3 2 + 1 = 3 3 = 3 (true)